THE THZ SPECTRUM OF LANTHANIDE AND TRANSITION METAL HALIDES - EFFECTS ON THE WATER SPECTRUM

\underline{G. SCHWAAB}, VINAY SHARMA AND MARTINA HAVENITH, Physical Chemistry II, Department of Chemistry and Biochemistry, Ruhr University Bochum, D-44780 Bochum, Germany.

Recently, we were able to show that aqueous solutions of alkali and earth-alkali halides show strong resonances in the THz range that may be attributed to movements of individual ions in their surrounding water cage. Here, we extend this study to bromides and chlorides of selected lanthanides and transition metals. The effects of these strong ions on the water network as well as the detectability of contact ion complexes will be discussed.